6) For each of the following sequences, write the next 3 terms. If arithmetic or
geometric, find t(n).
a) 4, 9, 14, 19, … b) 24, 12, 6, 3, ….
c) 91, 82, 73, 64, ….. d) 1, 4, 9, 16, …..
e) -9.5, -7, -4.5, -2, … f) -4, 8, -16, 32, ….
7) Find t(100) for all the arithmetic sequences from problem 6.

Answer by ewatrrr(24785)

a) 4, 9, 14, 19,24,29,34... Arithmetic sequence: d = 5, t(n) = 4 + 5(n-1) simplify.. 5n-1
b) 24, 12, 6, 3, 3/2, 3/4,3/8... . Geometric sequence: r = 1/2, t(n) = 24(1/2)^(n-1)
c) 91, 82, 73, 64, ….. Arithmetic sequence: d = -9, t(n) = 91 -9(n-1)= 100-9n
d) 1, 4, 9, 16, ….. t(n) = n^2
e) -9.5, -7, -4.5, -2, … Arithmetic sequence: d = 2.5, t(n) = -9.5 +2.5(n-1) = 2.5n -12
f) -4, 8, -16, 32, …. Geometric sequence: r = -1/2, t(n) = -4(-1/2)^(n-1)
7) Find t(100) for all the arithmetic sequences from problem 6.
plug and play
t(n) = 5n-1
t(n) = 24(1/2)^(n-1)
t(n) = 100-9n
t(n) = n^2
t(n) = 2.5n -12
t(n) = -4(-1/2)^(n-1)